# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 2.1 Lower Bound
Let $f_{1},\ldots,f_{m}\in\mathcal{F}$ be a collection be $\varepsilon$-separated functions, i.e., satisfying $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}^{2}>\varepsilon$ for each $i\neq j$. Suppose $J$ is uniformly distributed over the index set $[m]$, and $(Y_{i}|X_{i},J=j)=f_{j}(X_{i})+\xi_{i}$ where $\xi_{i}\sim N(0,\sigma^{2})$ for each $i\in[n],j\in[m]$. Then
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq\frac{\varepsilon^{2}}{4}\bigg{(}1-\frac{I(\vec{X},\vec{Y};J)+\log 2}{\log m}\bigg{)},$
@@
The next lemma gives an upper bound on the the mutual information term that appears in Fano’s Inequality.

### v5

@@ 2 Lower Bound
Let $f_{1},\ldots,f_{m}\in\mathcal{F}$ be a collection of $\varepsilon$-separated functions, i.e., satisfying $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\varepsilon$ for each $i\neq j$. Suppose $J$ is uniformly distributed over the index set $[m]$, and $(Y_{i}|X_{i},J=j)=f_{j}(X_{i})+\xi_{i}$ where $\xi_{i}\sim\mathcal{N}(0,\sigma^{2})$ for each $i\in[n],j\in[m]$. Then
$\inf_{\hat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{\vec{X},\vec{Y}}\|\hat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq\frac{\varepsilon^{2}}{4}\bigg{(}1-\frac{I(\vec{X},\vec{Y};J)+\log 2}{\log m}\bigg{)},$
@@
The next lemma gives an upper bound on the mutual information term that appears in Fano’s Inequality.